# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 5.3 Queue-based Methods are the worst ADL policies for the top winning trade
One result that we see empirically in §9 is that the Queue mechanism is by far the worst ADL policy for the top winning trade. We briefly formalize this with an informal proposition that we prove in Appendix B.3.
@@
Among all budget-balanced ADL policies with fixed severity $\theta$, the Queue policy (haircut largest winners first) uniquely minimizes the survival of the top winner. Let $\omega_{T}^{\text{Queue}}$ and $\omega_{T}^{\text{PR}}$ be the top winner’s post-ADL equity under Queue and Pro-Rata, respectively. Whenever the haircut budget $H=\theta D_{T}$ satisfies $H\leq e_{(1)}$, we have the strict gap
$\omega_{T}^{\text{PR}}-\omega_{T}^{\text{Queue}}\;=\;H\left(1-\frac{e_{(1)}}{W_{T}}\right)\;>\;0.$
@@
$\mathsf{PTSR}_{T}(\text{PR})-\mathsf{PTSR}_{T}(\text{Queue})\;=\;\theta\left(1-\frac{e_{(1)}}{W_{T}}\right).$
@@
The full proof is in Appendix B.3. Appendix B.5 further strengthens this statement by proving that Queue also minimizes the VaR and ES of the top survivor at every confidence level, so even classical risk metrics rank Queue as the worst policy for the top winner.
@@ 6 Fairness
The capped pro-rata rule is a water-filling algorithm that enforces per-account constraints before applying a pro-rata haircut. Combining haircut constraints (22) and equity constraints (23), we define an effective cap $\beta_{i}=\min\{\overline{h}_{i},1-\underline{e}_{i}/e_{T}(\mathfrak{p}_{i})\}$. The maximum haircut capacity is
$C(\beta)=\sum_{i=1}^{n}e_{T}(\mathfrak{p}_{i})\beta_{i}.$ (28)
@@
$\sum_{i=1}^{n}h_{\pi_{CP},i}e_{T}(\mathfrak{p}_{i})_{+}=\theta_{\pi}D_{T}(\mathcal{P}_{n}).$ (30)
This can be computed efficiently by sorting the caps $\beta_{i}$ and finding the threshold $\eta$ (see eq. (30) and the algorithm in Appendix D).
@@
We define three natural fairness properties:
  • Sybil resistance: Outcomes are identity-agnostic and invariant to splitting or merging accounts.
@@
  • Monotonicity (Stability): The policy preserves the relative ordering of post-haircut equity. If $e_{i}\geq e_{j}$, then $(1-h_{i})e_{i}\geq(1-h_{j})e_{j}$.
In Appendix C, we formalize the following result:
@@
This establishes capped pro-rata as the unique fair ordering rule. In contrast, queue-based rules (ranking by leverage or PNL) are generally not stable or sybil-resistant. RAP introduces a risk tilt, deviating from strict equity stability to reduce risk, but remains “as stable as possible” within its risk-weighted framework. In the numerical examples we present in the next section, we will see explicit examples where capped-pro rata is better for both exchange solvency and trader profitability than the PNL-leverage ranking and pure pro-rata.
@@ 7 Risk-aware Policies (RAP)
Fairness-focused rules like capped pro-rata treat all winning dollars equally, ignoring the heterogeneous risks imposed by different winners. Higher effective leverage implies thinner liquidation windows and greater sensitivity to execution costs [AlmgrenChriss2001, Gatheral2010]. We introduce Risk-Aware Pro-Rata (RAP), a family of rules that preserves the fairness structure of capped pro-rata while tilting haircuts toward higher-risk positions. This family includes Levered Pro-Rata (LPR) when the risk weighting is linear in leverage. We derive RAP from a robustness criterion that minimizes a one-step excess-deficit proxy.
@@
Appendix E.1 illustrates how pro-rata fails to control this deficit when shocks scale with leverage.
@@
Appendix E.1 provides numerical examples of these weightings.
@@ 7.1 Risk-Aware Pro-Rata (RAP) Rule
The RAP rule allocates haircuts using weights $w_{i}=\lambda_{T,i}g(\lambda_{T,i})$. With per-account caps $\beta_{i}$, the haircuts are:
$h_{\mathrm{RAP}(g),i}=\min\{\beta_{i},\tau w_{i}\},$ (33)
where $\tau$ is calibrated to meet the target budget. This concentrates haircuts on high-risk positions. See Appendix D for precise implementation of the algorithm.
@@ 7.2 Convex Dominance of RAP
We can construct an optimal risk model $g^{*}$ that minimizes $\delta_{T}$. Intuitively, an optimal risk model should maximize the marginal reduction of $\delta_{T}$ per unit of equity $e_{T,i}$. Differentiating $\delta_{T}$ with respect to the haircut $h_{T,i}$ reveals:
$\frac{\partial\,\delta_{T}(h)}{\partial h_{T,i}}\ =\ n_{T,i}\,\psi_{i}\!\left(\tfrac{1}{\lambda_{T,i}}\right)\ =\ e_{T,i}\,\lambda_{T,i}\,\psi_{i}\!\left(\tfrac{1}{\lambda_{T,i}}\right)\ =\ e_{T,i}\,\rho_{i}(\lambda_{T,i}),$ (34)
where we define the perspective transform $\rho_{i}(\lambda)=\lambda\,\psi_{i}(1/\lambda)$. The greedy rule that maximizes the marginal reduction per equity-dollar is exactly the rule that prioritizes winners with the largest $\rho_{i}(\lambda_{T,i})$.
@@
The function $\rho_{i}$ is the perspective transform of the convex function $\psi_{i}$ [BoydVandenberghe2004, §2.3.3], widely used to induce 1-homogeneity. Economically, $u=1/\lambda$ is the bankruptcy buffer, which is the percentage price shock required to wipe out the position’s equity (since loss equals equity when shock magnitude $Z=e/\tilde{n}=1/\lambda$). Using convex duality, we can write $\rho_{i}(\lambda)=\sup_{y\geq 0}\{y-\lambda\,\psi_{i}^{*}(y)\}$, where $y$ is a ‘yield’ (equity change per unit buffer change) and $\psi_{i}^{*}(y)$ is the cost to insure a position with yield $y$. Thus, maximizing $\rho_{i}$ is equivalent to maximizing the net solvency gain per unit of equity. For example, if a trader with leverage $\lambda=10$ faces an insurance cost of 0.5% for a yield of 8%, the net solvency gain is 0.0810(0.005)=0.030.08-10(0.005)=0.03 per unit equity.
@@
The proof is in Appendix E.2. This result implies that if we have prior knowledge about the kernel $K$ (e.g., from historical data or backtesting), then constructing $g^{*}$ via the perspective transform yields the optimally robust and fair haircut model.
@@ 7.3 Limitations of One-Step Optimality
Consider sequential shocks $Z_{t+1},Z_{t+2}$ with positive correlation. High-leverage positions might act as a hedge against future shocks. In such cases, aggressively liquidating high-leverage winners (as in RAP) can increase the two-step deficit compared to pro-rata. Appendix E.1 details an AR(1) shock example where the sum of deficits is lower for pro-rata.
@@
Exchanges must also consider exchange long–term value (LTV)—the future fees and liquidity provided by traders. RAP targets high-leverage, often high-volume, traders. Fully liquidating these accounts to minimize immediate risk can reduce the exchange’s long-term utility. Appendix E.3 provides an example where preserving a high-value trader (via pro-rata) yields higher total utility than the risk-minimizing RAP.
@@ 8 Multiple–Round ADL as a Stackelberg Control Problem
As the previous section illustrates, risk-aware pro-rata mechanisms work well as robust, static strategies but not in a dynamic setting. Both exchange future revenue and correlated shocks require including a model of the dynamics of ADL over multiple rounds. We use the intuition gained from RAP mechanisms for a single round to help extend the model to the multiple round setting. A common place where there is a difference between one-shot equilibria and multi-round equilibria is in the difference between Nash and Stackelberg games. We will illustrate how the multi-round ADL problem is naturally a Stackelberg game whereas the ADL policies that we have studied so far are more like Nash optimal strategies. To do this, we will first briefly describe Stackelberg games and then provide examples demonstrating that it is the right formalism for this extension.
@@ 8.1 ADL is a Stackelberg Game
A Stackelberg (leader–follower) game is a sequential game where the leader first commits to a strategy or policy, the follower observes this commitment, and then best–responds; the leader chooses its commitment anticipating this response [Stackelberg1934, FudenbergTirole1991, BasarOlsder1999]. [^11] In the finance and control literature, this setup is often framed as “Dynamic Optimal Control with Rational Expectations” or a principal-agent problem with commitment [Puterman1994, Bertsekas2017, LaffontMartimort2002]. The defining features are (i) observable commitment, (ii) sequential move order, and (iii) the leader’s optimization against the follower’s best–response correspondence. In repeated or multi–round settings, the leader can learn a near–optimal commitment by updating its policy from round to round using feedback from realized follower responses and shocks; this supports standard online learning guarantees when losses are convex (or admit convex surrogates) [Zinkevich2003, ShalevShwartz2012, Hazan2019], and, under additional regularity, learnability results specific to Stackelberg games [conitzer2006].
@@
Credibility and Time Inconsistency.
Crucially, the Stackelberg formulation assumes the leader can credibly commit to a policy path. In decentralized venues (e.g., Hyperliquid), this is enforced via on-chain code or DAO governance, making the commitment verifiable. In centralized venues (e.g., Binance), the commitment is reputational. Without commitment, the game reverts to a Nash setting where the exchange suffers from time inconsistency: ex-ante, it prefers a “gentle” policy to attract liquidity, but ex-post (once a deficit $D_{t}$ materializes), it has a dominant incentive to defect to maximal severity (Queue ADL) to minimize immediate insolvency risk. This temptation to defect undermines liquidity provision unless the commitment is binding, further justifying the Stackelberg framework over pure control theory.
@@
Nash equilibria are one–shot, simultaneous–move fixed points where each player best–responds to the other’s action but no one commits ex ante [FudenbergTirole1991]. By contrast, Stackelberg equilibria embed the leader’s commitment advantage, often producing different (and more robust or revenue–enhancing) outcomes [Stackelberg1934, BasarOlsder1999]. Static ADL strategies such as venue queues (Binance/Hyperliquid) or single–round RAP can be analyzed under a Nash lens, as we assume there is no further response by a leader after a single execution of an ADL policy. In practice, ADL controllers are used repeatedly (such as on October 10, 2025; see §9) which more resembles a Stackelberg game. In Appendix F.3 we give an explicit example for ADL where a Stackelberg equilibrium differs from a Nash outcome and in which the Stackelberg equilibrium strictly improves exchange revenue and robustness.
Robustness and Learnability of Stackelberg games.
Robust Stackelberg models capture the leader’s desire to hedge against uncertainty hidden in the follower’s strategy by optimizing for worst–case responses [aghassi2006]. In ADL, robustness corresponds to resilience against adverse price/oracle shocks, liquidity droughts, and funding–rate shocks. Specifically, the exchange chooses a policy that anticipates such perturbations and steers the system toward states with smaller tail externalities (analogous to Stackelberg Security Games [Tambe2011, conitzer2006]).
When the leader’s per round objective function is convex in its policy (or admits convex surrogates) and constraints are convex, standard online methods (mirror descent, FTRL) achieve sublinear regret and convergence to near–optimal commitments [Zinkevich2003, ShalevShwartz2012, Hazan2019]. Recent work studies gradient–based learning in Stackelberg games directly, providing conditions under which leader updates converge to Stackelberg equilibria, including bandit/noisy feedback regimes. These tools align well with ADL’s multi–round nature and with the RAP geometry used for choosing haircuts.
@@ 8.2 Why do we need a dynamic model?
To compare policies dynamically, we track two opposing clocks starting from a default event at $\tau_{\mathrm{def}}$:
@@
Intuitively, $\tau_{\mathrm{solv}}$ measures how fast the “hole is plugged,” while $\tau_{\mathrm{rev}}$ measures how fast the “business recovers.” Static policies face a fundamental trade-off between these times, governed by the concentration of haircuts.
@@
We can rigorously rank policies using the Schur-convex (majorization) order from §3. Consider two canonical policies:
  • Queue ADL (Concentrated): This policy fully haircuts the largest winner until the deficit is covered. It is extremely Schur-convex (unequal). By targeting the deepest pocket, it clears the deficit with the minimal total haircut volume, minimizing $\tau_{\mathrm{solv}}$. However, it destroys the highest-value traders (the “whales”), causing a massive drop in LTV and maximizing $\tau_{\mathrm{rev}}$.
  • Pro-Rata ADL (Diffused): This policy spreads the haircut proportionally across all winners. It is Schur-concave (equal). It preserves the whales but requires a larger total haircut volume to clear the same deficit (due to many small accounts having insufficient equity), effectively delaying $\tau_{\mathrm{solv}}$. However, because the per-trader impact is small, churn is low, and $\tau_{\mathrm{rev}}$ is minimized.
This intuition is formalized in the following proposition, which states that any policy that concentrates haircuts (like Queue) will recover solvency faster but revenue slower than a diffuse policy (like Pro-Rata).
@@
Under separable convex stage losses and monotone shock dynamics, if the residual exposure of policy $A$ majorizes policy $B$ (e.g., Queue vs. Pro-Rata), then:
@@
Endogenous Price Feedback.
Static models typically treat the deficit $D_{t}$ as exogenous. In reality, large ADL events create endogenous feedback: the act of closing positions (even if socialized among winners) reduces open interest and can signal distress, causing market makers to widen spreads or withdraw bids. This secondary liquidity shock can drive mark prices further against the exchange, increasing $D_{t+1}$. Dynamic policies like $\pi_{\text{md}}$8.5) naturally handle this by treating the environment as adversarial (or responsive), whereas static policies risk under-estimating the “death spiral” effect of the ADL action itself.

### v3

@@ 5.3 Queue-based Methods are the worst ADL policies for the top winning trade
One implication of the theory is that the Queue mechanism is by far the worst ADL policy for the top winning trade. We briefly formalize this with an informal proposition that we prove in Appendix C.3.
@@
Among all valid ADL policies with fixed severity $\theta$, the Queue policy (haircut largest winners first) uniquely minimizes the survival of the top winner’s endowment. Let $\upsilon_{T}^{\text{Queue}}$ and $\upsilon_{T}^{\text{PR}}$ be the top winner’s post-ADL endowment under Queue and Pro-Rata, respectively. Whenever the haircut budget $H=\theta D_{T}$ satisfies $H\leq w_{(1)}$ (where $w_{(1)}$ is the largest endowment), we have the strict gap
$\upsilon_{T}^{\text{PR}}-\upsilon_{T}^{\text{Queue}}\;=\;H\left(1-\frac{w_{(1)}}{U_{T}}\right)\;>\;0.$
@@
$\mathsf{PTSR}_{T}(\text{PR})-\mathsf{PTSR}_{T}(\text{Queue})\;=\;\theta\left(1-\frac{w_{(1)}}{U_{T}}\right).$
@@
The full proof is in Appendix C.3. Appendix C.5 further strengthens this statement by proving that Queue also minimizes the VaR and ES of the top survivor’s endowment at every confidence level, so even classical risk metrics rank Queue as the worst policy for the top winner.
@@ 6 Fairness
The capped pro-rata rule is a water-filling algorithm that enforces per-account constraints before applying a pro-rata haircut. Combining haircut constraints (26) and equity constraints (27), we define an effective cap $\beta_{i}=\min\{\overline{h}_{i},1-(\underline{e}_{i}-c_{i,T})/w_{i,T}\}$ for accounts with $w_{i,T}>0$, where the equity constraint is reinterpreted in terms of the minimum post-ADL endowment. This $\beta_{i}$ is the most we can haircut account $i$’s endowment while honoring both its per-account haircut limit and its minimum post-ADL equity. The maximum haircut capacity is
$C(\beta)=\sum_{i=1}^{n}w_{i,T}\beta_{i}.$ (32)
@@
$\sum_{i=1}^{n}h_{\pi_{CP},i}w_{i,T}=\theta_{\pi}D_{T}(\mathcal{P}_{n}).$ (34)
This can be computed efficiently by sorting the caps $\beta_{i}$ and finding the threshold $\eta$ (see eq. (34) and the algorithm in Appendix E).
@@
Fix round $T$ and policy $\pi$. Let $x_{\pi,i}=h_{\pi,i}w_{i,T}$ denote seized endowment and $s_{\pi,i}=w_{i,T}-x_{\pi,i}$ denote surviving endowment. We define three natural fairness properties:
  • Sybil resistance: For any split $w_{i,T}=\sum_{a}z_{a}$, with $\tilde{x}_{\pi,a}$ denoting seizure on split children, aggregate seizure is invariant: $\sum_{a}\tilde{x}_{\pi,a}=x_{\pi,i}$.
@@
  • Monotonicity (Stability): The policy preserves endowment ordering: if $w_{i,T}\geq w_{j,T}$, then $s_{\pi,i}\geq s_{\pi,j}$ (and consequently, if $e_{i}\geq e_{j}$ pre-ADL and cash components are comparable, then $e^{\prime}_{i}\geq e^{\prime}_{j}$ post-ADL).
In Appendix D, we formalize the following result:
@@
This establishes capped pro-rata as the unique fair ordering rule. For queue-based rules, Sybil resistance depends on the ranking score (Appendix D, Proposition D.1): absolute-value scores (e.g. total PNL) are Sybil vulnerable, while ratio scores are Sybil resistant when proportional splitting does not create score ties. Hyperliquid’s documented ADL score
$\frac{\mathrm{mark}}{\mathrm{entry}}\cdot\frac{\mathrm{notional}}{\mathrm{account\ value}}$
is degree-zero in $(\mathrm{notional},\mathrm{account\ value})$ and therefore split-invariant under proportional scaling [72]. [^11] Worked example (also discussed in [1]): mark/entry =1.2=1.2, notional/account value =12,000,000/2,500,000=4.8=12{,}000{,}000/2{,}500{,}000=4.8, so score $5.76$. Splitting into $100$ accounts gives 120,000/25,000=4.8120{,}000/25{,}000=4.8 per account, so each score remains $5.76$. However, Sybil resistance does not imply monotonicity: Appendix D, Proposition D.3 gives a two-account counterexample where queue budget exhaustion reverses post-ADL ordering. RAP introduces a risk tilt, deviating from strict equity stability to reduce risk, but remains “as stable as possible” within its risk-weighted framework. In the numerical examples we present in the next section, we will see explicit examples where capped pro-rata is better for both exchange solvency and trader profitability than the Queue and Pro-Rata.
@@ 7 Risk-aware Policies (RAP)
Fairness-focused rules like capped pro-rata treat all winning dollars equally, ignoring the heterogeneous risks imposed by different winners. Higher effective leverage implies thinner liquidation windows and greater sensitivity to execution costs [8, 58]. We introduce Risk-Aware Pro-Rata (RAP) rules, a family of rules that preserves the fairness structure of capped pro-rata while tilting haircuts toward higher-risk positions. This family includes Levered Pro-Rata (LPR) when the risk weighting is linear in leverage. We derive RAP from a robustness criterion that minimizes a one-step excess-deficit proxy.
@@
Appendix F.1 illustrates how pro-rata fails to control this deficit when shocks scale with leverage.
@@
Appendix F.1 provides numerical examples of these weightings.
@@ 7.1 Risk-Aware Pro-Rata (RAP) Rule
The RAP rule allocates haircuts using risk weights $\tilde{w}_{i}=\lambda_{T,i}g(\lambda_{T,i})$ (not to be confused with the endowment $w_{i,T}$). With per-account caps $\beta_{i}$, the haircuts on the endowment are:
$h_{\mathrm{RAP}(g),i}=\min\{\beta_{i},\tau\tilde{w}_{i}\},$ (37)
where $\tau$ is calibrated to meet the target budget $\sum_{i}h_{\mathrm{RAP}(g),i}w_{i,T}=\theta D_{T}$. This concentrates haircuts on high-risk positions (those with high leverage and high risk scores). See Appendix E for precise implementation of the algorithm.
@@ 7.2 Convex Dominance of RAP
We can construct an optimal risk model $g^{*}$ that minimizes $\delta_{T}$. Intuitively, an optimal risk model should maximize the marginal reduction of $\delta_{T}$ per unit of endowment $w_{T,i}$ seized. Differentiating $\delta_{T}$ with respect to the seized amount $x_{T,i}=h_{T,i}w_{T,i}$ reveals:
$\frac{\partial\,\delta_{T}(x)}{\partial x_{T,i}}\ =\ n_{T,i}\,\psi_{i}\!\left(\tfrac{1}{\lambda_{T,i}}\right)\ =\ e_{T,i}\,\lambda_{T,i}\,\psi_{i}\!\left(\tfrac{1}{\lambda_{T,i}}\right)\ =\ e_{T,i}\,\rho_{i}(\lambda_{T,i}),$ (38)
where we define the perspective transform $\rho_{i}(\lambda)=\lambda\,\psi_{i}(1/\lambda)$. The greedy rule that maximizes the marginal reduction per endowment-dollar seized is exactly the rule that prioritizes winners with the largest $\rho_{i}(\lambda_{T,i})$. Note that while the gradient is expressed in terms of equity $e_{T,i}$ (since insolvency is an equity concept), the control variable is the endowment $w_{T,i}$.
@@
The function $\rho_{i}$ is the perspective transform of the convex function $\psi_{i}$ [29, §2.3.3], widely used to induce 1-homogeneity. Economically, $u=1/\lambda$ is the bankruptcy buffer, which is the percentage price shock required to wipe out the position’s equity (since loss equals equity when shock magnitude $Z=e/\tilde{n}=1/\lambda$). Using convex duality, we can write $\rho_{i}(\lambda)=\sup_{y\geq 0}\{y-\lambda\,\psi_{i}^{*}(y)\}$, where $y$ is a ‘yield’ (equity change per unit buffer change) and $\psi_{i}^{*}(y)$ is the cost to insure a position with yield $y$. Thus, maximizing $\rho_{i}$ is equivalent to maximizing the net solvency gain per unit of endowment seized (the gradient in (38) is expressed in equity terms because insolvency is an equity concept, but the control variable is the endowment $w_{T,i}$). For example, if a trader with leverage $\lambda=10$ faces an insurance cost of 0.5% for a yield of 8%, the net solvency gain is 0.0810(0.005)=0.030.08-10(0.005)=0.03 per unit equity, which translates to a gain per unit of endowment seized under the PNL-only model.
@@
The proof is in Appendix F.2. This result implies that if we have prior knowledge about the kernel $K$ (e.g. from historical data or backtesting), then constructing $g^{*}$ via the perspective transform yields the optimally robust and fair haircut model.
@@ 7.3 Limitations of One-Step Optimality
Consider sequential shocks $Z_{t+1},Z_{t+2}$ with positive correlation. High-leverage positions might act as a hedge against future shocks. In such cases, aggressively liquidating high-leverage winners (as in RAP) can increase the two-step deficit compared to pro-rata. Appendix F.1 details an AR(1) shock example where the sum of deficits is lower for pro-rata.
@@
Exchanges must also consider exchange long–term value (LTV) — the future fees and liquidity provided by traders. RAP targets high-leverage, often high-volume, traders. Fully liquidating these accounts to minimize immediate risk can reduce the exchange’s long-term utility. Appendix F.3 provides an example where preserving a high-value trader (via pro-rata) yields higher total utility than the risk-minimizing RAP.
@@ 8 Multiple–Round ADL as a Stackelberg Control Problem
As the previous section illustrates, risk-aware pro-rata mechanisms work well as robust, static strategies but not in a dynamic setting. Both future exchange revenue and correlated shocks require including a model of the dynamics of ADL over multiple rounds. We use the intuition gained from RAP mechanisms for a single round to help extend the model to the multiple round setting. A common place where there is a difference between one-shot equilibria and multi-round equilibria is in the difference between Nash and Stackelberg games. We will illustrate how the multi-round ADL problem is naturally a Stackelberg game whereas the ADL policies that we have studied so far are more like Nash optimal strategies. To do this, we will first briefly describe Stackelberg games and then provide examples demonstrating that it is the right formalism for this extension.
@@ 8.1 ADL is a Stackelberg Game
A Stackelberg (leader–follower) game is sequential: the leader first commits to a policy, the follower observes it, and then best–responds; the leader chooses its commitment anticipating that response [117, 57, 18]. [^12] In finance and control, this appears as “dynamic optimal control with commitment” or a principal–agent model with rational expectations [97, 20, 80]. Commitment here means the policy is observable and credible to followers. The move order is therefore leader first, follower second, and the leader optimizes against the follower’s best–response set. Over many rounds, the leader can learn a near–optimal commitment by updating $\{\theta_{t},h_{t}\}$ using feedback from realized follower responses and shocks, recovering standard online-learning guarantees when losses are convex (or have convex surrogates) [118, 103, 66, 43]. The Stackelberg view also supports robust design: the exchange can pick policies that hedge against worst–case follower responses and price/oracle shocks, steering the system toward states with smaller tail externalities [6, 113].
@@
Single vs. Multi-Round Models.
One-shot (static) ADL treats each deficit in isolation, optimizing a fixed severity $\theta$ and haircut rule $h$ for that round. The multi-round Stackelberg model links rounds: the leader commits to a policy path $\{\theta_{t},h_{t}\}$, and followers adapt, creating state dependence via churn, liquidity, and leverage. This coupling makes commitment valuable: a policy that is optimal in one round can be dominated once cross-shock feedback is considered.
@@
Nash equilibria are one–shot, simultaneous–move fixed points where each player best–responds to the other’s action but no one commits ex ante [57]. By contrast, Stackelberg equilibria embed the leader’s commitment advantage, often producing different (and more robust or revenue–enhancing) outcomes [117, 18]. Static ADL strategies such as venue queues (Binance/Hyperliquid) or single–round RAP can be analyzed under a Nash lens, as we assume there is no further response by a leader after a single execution of an ADL policy. In practice, ADL controllers are used repeatedly (such as on October 10, 2025; see §9) which more resembles a Stackelberg game. In Appendix G.3 we give an explicit example for ADL where a Stackelberg equilibrium differs from a Nash outcome and in which the Stackelberg equilibrium strictly improves exchange revenue and robustness.
@@ 8.2 Why do we need a dynamic model?
To compare policies dynamically, we track two opposing clocks starting from a default event at $\tau_{\mathrm{def}}$: $\tau_{\mathrm{solv}},\tau_{\mathrm{rev}}$. Intuitively, $\tau_{\mathrm{solv}}$ is how quickly deficits are cleared whereas $\tau_{\mathrm{rev}}$ is how quickly fee revenue rebounds.
@@
Formally,
$\tau_{\mathrm{solv}}=\inf\{t\geq\tau_{\mathrm{def}}:\mathsf{IF}_{t}\geq\delta\},\qquad\tau_{\mathrm{rev}}=\inf\{t\geq\tau_{\mathrm{def}}:\mathop{\bf E{}}[\mathrm{LTV}_{t}]\geq(1-\epsilon)\mathrm{LTV}_{\text{pre}}\}.$
As we will show, static policies trade these quantities off one another off. We also show that the concentration of how haircuts are distributed amongst winning traders controls this trade-off.
@@
We can rank policies by how concentrated their haircuts are using the Schur order (§3):
  • Queue (concentrated): Haircuts hit the largest winners first, yielding a highly unequal (Schur-convex) vector. Deficits clear quickly (small $\tau_{\mathrm{solv}}$), but the biggest fee payers churn, so LTV recovers slowly (large $\tau_{\mathrm{rev}}$).
  • Pro-Rata (diffuse): Haircuts are spread evenly (Schur-concave). Solvency takes longer because shallow accounts contribute little (larger $\tau_{\mathrm{solv}}$), but everyone stays engaged, so revenue rebounds faster (small $\tau_{\mathrm{rev}}$).
Any intermediate policy traces this frontier: more concentration shortens $\tau_{\mathrm{solv}}$ but lengthens $\tau_{\mathrm{rev}}$. This intuition is formalized in the following proposition, which states that any policy that concentrates haircuts (like Queue) will recover solvency faster but revenue slower than a diffuse policy (like Pro-Rata).
@@
Under separable convex stage losses and monotone shock dynamics, if the residual exposure of policy $A$ majorizes policy $B$ (e.g. Queue vs. Pro-Rata), then: